OpenStudy (anonymous):

A man 5'8" wishes to find the height of an oak tree in his front yard.He walks along the shadow of the tree untill his head is in a position where the end of his shadow exactly overlap the end ofthe treetop's shadow.He is now 11'3" from the foot of the tree an 8'6" from the end of the shadows.How tall is the oak tree?

OpenStudy (anonymous):

if you take height of man as opposite side of an angle and length of his shadow as adjacent side of that angle then tan(x)=5'8"/8'6"=68"/102" same in the case of tree . let h be height of tree. tan(x)=h/(11'3"+8'6")=68/102. h/237=68/102, h=158"=13'2"